OpenStudy (anonymous):

Describe the difference between conservative and non-conservative forces as it pertains to work.

OpenStudy (anonymous):

Conservative forces have Triangle.W = 0

OpenStudy (anonymous):

Because they follow a closed path

OpenStudy (anonymous):

Conservative forces do the same amount of work going from A to B regardless of the path, like lifting against gravity.

OpenStudy (anonymous):

what about non-conservative forces ....

OpenStudy (anonymous):

Think for yourself.. think about friction.. imagine you move a block from one corner in your bedroom, to another corner do you think the amount of work done depends on what path you take? or not?

OpenStudy (anonymous):

Sorry, thought that would be clear: they do depend on the path taken.

OpenStudy (anonymous):

a force is said to be conservative if the work done by the force (or against the force) in moving a body depends only upon the initial and the final positions of the body and is independent of the path followed between the initial and the final positions. (in this fig. consider alpha to be theta..it was easier to draw alpha ....so..i drew alpha instead....but consider it to be theta) in the 1st case the object is simply being taken upto a height(h) without imparting any velocity.. work done= force * displacement = mg*h = mgh in 2nd case the object is taken along a smooth inclined plane of height h and inclination theta.the force =mgsin(theta) without imparting any velocity to the object.but the dispacement = slope.=h/sin(theta)  work =force * displacement =mgsin(theta) * h/sin(theta) =mgh in 3rd case each vertical path is of height (h0) and the height of the topmost point is h. so work done =\[\Sigma mg \times (h)= mgh \] non conservative force:- a force is said to be non conservative if the work done by the force (or against the force) in moving a body depends upon the path followed between the two points. eg. friction. suppose a car has to go from point A to Point B. it has 2 paths to reach B.One is path made of Ice and the other is of concrete....the work done by friction is diff. in both the cases as the friction coefficient is different in both the cases..